OpenStudy (anonymous):

Calc Q: The owner of a play equipment store wants to construct a fence to enclose a rectangular outdoor storage area by using part of one side of the store, which is 250 feet long, as one of the sides of the enclosed area. There are 470 feet of fencing available for the other three sides of the enclosure. Find the dimensions of the outdoor enclosure with the most area. Note:

OpenStudy (anonymous):

You can do this, no problem just think of two key formulas, the area and the permiter

OpenStudy (anonymous):

470=L+2w, only one L cause we are dealing with only 3 sides

OpenStudy (anonymous):

then you know that Area=LW, but from the formula above, we can rearange to get L, right? plug that into the formula for the area and bam, differentiate that.

OpenStudy (anonymous):

470=L+2w------>L=?

OpenStudy (anonymous):

(470-L)/2 right?

OpenStudy (anonymous):

not really :) we want to find L, what does L equal?

OpenStudy (anonymous):

oh ok ma bad... its L=2W-450

OpenStudy (anonymous):

you mean 470

OpenStudy (anonymous):

yes sorry hehe

OpenStudy (anonymous):

so what about the 250? Do i need to work with it?

OpenStudy (anonymous):

naw, its a not useful in this case

OpenStudy (anonymous):

oh ok...

OpenStudy (anonymous):

but make sure you do the differentiation okay

OpenStudy (anonymous):

yes...so after I did the diff and then what?

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

W=235

OpenStudy (anonymous):

okay, then all you have to do is find out what L is

OpenStudy (anonymous):

By doing this d/dx (2w-470) :)

OpenStudy (anonymous):

L=0?

OpenStudy (anonymous):

no way jose :)

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

sory cruzermanth my connection was bad...i missed what u were saying earlier